OpenStudy (anonymous):

What is the equation, in point-slope form, of the line that has a slope of 6 and passes through the point (–1, –8)? mc002-1.jpg mc002-2.jpg mc002-3.jpg mc002-4.jpg

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) slope(m) = 6 (-1,-8)...x1 = -1 and y1 = -8 now sub, but do not solve...leave the parenthesis in y - (-8) = 6(x - (-8) y + 8 = 6(x + 8) <== point slope form

OpenStudy (anonymous):

the answer choices is y+8=6(x+1) y-8=6(x-1) y+1=6(x+8) y-1=6(x-8)

OpenStudy (texaschic101):

oops...sorry...I subbed wrong It should be : y + 8 = 6(x + 1)

OpenStudy (anonymous):

thankss